Consider a town in which only two residents, Jacques and Kyoko, own wells that produce water safe for drinking. Jacques and Kyoko can pump and sell as much water as they want at no cost. For them, total revenue equals profit. The following table shows the town's demand schedule for water. Price Quantity Demanded Total Revenue (Dollars per litre) Litres of water) (Dollars,) 4.20 3.85 3.50 3.15 2.80 2.45 2.10 1.75 1.40 1.05 0.70 0.35 40 80 120 160 200 240 280 320 360 400 440 480 $154.00 $280.00 $378.00 $448.00 $490.00 $504.00 $490.00 $448.00 $378.00 $280.00 $154.00 Suppose Jacques and Kyoko form a cartel and behave as a monopolist. The profit-maximizing price is $ is per litre, and the total output litres. As part of their cartel agreement, Jacques and Kyoko agree to split production equally. Therefore, Jacques's profit is and Kyoko's profit is $ Suppose that Jacques and Kyoko have been successfully operating as a cartel. They each charge the monopoly price and sell half of the monopoly quantity. Then one night before going to sleep, Jacques says to himself, "Kyoko and I aren't the best of friends anyway. If I increase my production to 40 litres more than the cartel amount, I can increase my profit even though her profit goes down. I will do that starting tomorrow." to After Jacques implements his new plan, the price of water levels, Jacques's profit becomes $ per litre. Given Kyoko and Jacques's production and Kyoko's profit becomes $
Because Jacques has deviated from the cartel agreement and increased his output of water to 40 litres more than the cartel amount, Kyoko decides that she will also increase her production to 40 litres more than the cartel amount. After Kyoko increases her production, Jacques's profit becomes S sum of the profits of Jacques and Kyoko) is now $ ,Kyoko's profit becomes S and total profit (the True or False: Based on the fact that both Jacques and Kyoko increased production from the initial cartel quantity, you know that the output effect wa arger than the price effect at that quantity. True O False Jacques and Kyoko have each cheated on their cartel agreement and increased production by 40 litres more than the cartel amount. However, they both realize that if they continue to increase output beyond this amount, they'll each suffer a decrease in profit. (To see this for yourself, consider Jacques's profit when he produces 80 litres more than the cartel amount compared to his profit when he produces 40 litres more than the cartel amount.) Neither Jacques nor Kyoko has an incentive to further increase output, nor does either have an incentive to decrease output. This outcome is an example of
Answer #1

When they form a cartel and behave as a monopolist,they will set price and quantity where the revenue is maximum, so the profit-maximizing price is 2.10 and the total output is 240.As part of their agreement,the production is split equally so Jacques profit is 504/2 = 252 and Kyoko profit is 504/2 = 252

When Jacques implements his new plan and increases his production by 40 more litres then price decreases to 1.75 so Jacques profit becomes 240/2+40 = 160*1.75 = 280 and Kyoko profit becomes 120*1.75 = 210

When Kyoko increases her production by 40 litres,then the price will further decrease to 1.40 so Jacques profit becomes 320/2= 160*1.40 = 224 and Kyoko profit becomes 160*1.40 =224 so total profit is now 224+224 = 448

You know that the output effect is larger than the price effect - True

This outcome is an example of a Nash Equilibrium